OpenStudy (anonymous):

please solve this trig question sin x/1-cos x + 1-cosx/sinx

OpenStudy (dan815):

can u clarify question by putting brackets wherever they need to go

OpenStudy (anonymous):

do you know how to solve basic trigonometric equations?

OpenStudy (anonymous):

or simpler do you know how to add fractions with different denominators

hero (hero):

\(\dfrac{\sin(x)}{1 - \cos(x)} + \dfrac{1 - \cos(x)}{\sin(x)}\)